2. Given P(E) = 0.25, P(F) = 0.6, and P(E ∪ F) = 0.7.

Find:


a. What is P(E ∩ F)?

b. Are event E and event F mutually exclusive? Justify your answer.

c. Are event E and event F independent? Justify your answer.


1
Expert's answer
2020-09-22T16:22:39-0400

"P(E \cup F) = P(E) + P(F) - P(E \cap F)". From here we have:

a. "P(E \cap F) = P(E) + P(F) - P(E \cup F) = 0.25+0.6-0.7 =0.15"

b. If events are mutually exclusive, then "P(E \cap F) =0". As we see from (a), "P(E \cap F) \neq 0", therefore these events are not mutually exclusive.

c. If events are independent, then "P(E \cap F) = P(E) \cdot P(F)". Let's check this.

"P(E) \cdot P(F) = 0.25 \cdot0.6 = 0.15" and "P(E \cap F)= 0.15".

Indeed, "P(E \cap F) = P(E) \cdot P(F)", so these events are independent.